# Vessel Detection and Localization Using Distributed Acoustic Sensing in Submarine Optical Fiber Cables

### v1

@@ V-B Spectral Feature Extraction
Extracting robust features from DAS data for vessel detection and localization is complicated due to the intricate physical processes converting ship noise into measured strain. Accurately modeling this conversion is challenging due to several factors: the complex and variable nature of noise radiated by different vessels [14]; the complexities of underwater acoustic propagation, including attenuation, reflection, and Doppler effects [25, 40, 41]; variations in the coupling between the fiber cable and seabed, which strongly influence strain measurements [42, 37]; and potential artifacts introduced by the DAS system itself (due to electronic and optical design, and the conditions in the interrogator deployment environment).
Given that acquiring the precise knowledge needed for a full physics-based model is often impractical due to environmental variability and incomplete deployment information, our proposal is adopting a purely data-driven strategy. A frequent technique involves analyzing the energy within specific frequency bands, which has proven effective for interpreting DAS signals and providing features for associated ML tasks [43, 44, 10]. Adopting this principle, our work utilizes a data-driven spectral analysis, informed by available AIS data, to generate feature vectors suitable for guiding the ML process. Further details are given in Section VI-C, as it implies experimental work for feature vector design.
@@ V-C Machine Learning Tasks and Training Modules
  • XGBoost (eXtreme Gradient Boosting): An optimized gradient boosting framework that learns an ensemble of decision trees [45]. XGBoost was selected for its proven effectiveness, scalability to large datasets, and inherent mechanisms to mitigate overfitting, making it suitable for handling potentially high-dimensional feature spaces derived from DAS data.
@@
Another key element that impacts the reliability and robustness of a system and its real generalization capabilities is the rigor of the experimental approach [46]. In time-series data, such as DAS recordings, temporal dependencies can introduce biases if not properly handled. The typically used random partitioning strategy may result in training and validation sets containing temporally adjacent data, leading to information leakage and overestimation of model performance. Special care must be taken in this stage, adopting specific cross-validation strategies that include restrictions on the selection of temporal ranges for training and testing subsets.
@@ V-D Result Post-Processing
In ML tasks dealing with time series, we can further improve the classification results by applying post-processing methods. Typical strategies are smoothing via sliding window averaging, median or mode filtering, majority voting (using ensembles or temporal ranges), multiple intervals, etc. [47].
Finally, one key benefit of DAS is the extensive range of sensed positions, which allows the systems to exploit the spatial and temporal redundancy of the acquired data. This should be explicitly included in either the feature extraction or ML stages. Suitable alternatives include exploiting array signal processing [30], or data integration across temporal and spatial ranges [48, 49], which will be detailed in Section VI-D.
@@ VI-A Database Partitioning
Given the characteristics of our task, we adopted the $k$-fold cross-validation strategy, a widely used technique for assessing model generalization capability [50]. This involves dividing the whole dataset into $k$ equal subsets (folds). The model is trained on k1k-1 folds and validated (tested) on the remaining fold. This process repeats $k$ times, with each fold serving once as the test set. The results from all test sets are then averaged to estimate the overall performance. Therefore, this method ensures that every data point is used for both training and testing, providing a comprehensive evaluation of the model’s performance while maintaining a strict separation between training and testing data.
@@ VI-B Performance Metrics
  • $F_{1}{}$-score: The harmonic mean of $Precision$ and $Recall$ (the higher, the better), computed as: $F_{1}{}\nobreak=\nobreak 2\cdot\nobreak\frac{Precision\cdot Recall}{Precision+Recall}$, where $Precision$ quantifies the number of correctly detected positive samples out of all the samples classified as positive, while $Recall$ quantifies the percentage of positive samples that are correctly detected.
@@
In the vessel distance-estimation task, where the MAE will be calculated on data subsets with different distance thresholds, we will also normalize the MAE by these distance thresholds ($\delta_{thr}\in D_{thr}$), to get a relative estimation of the error, as compared with this distance threshold: $rMAE=\nobreak\frac{MAE}{\delta_{thr}}$.
To compare the results between the two ML approaches, we will quantify the uncertainty of our performance metrics using a non-parametric bootstrap procedure [51] (adapted to our $k$-fold cross-validation strategy) that has been further validated by subsequent works on bootstrap methods in ML [52, 53].
@@ VI-C Feature Extraction Analysis
(a) Spectral content in the (0.01(0.01-250)Hz250)\,Hz range.
@@
Figure 7: Long term spectral averages for different environmental conditions.
Figure 8: Feature vector energy map across sensed fiber positions (top) vs. minimum distance to nearest vessel (bottom). Horizontal axis is time.
Fig. 7.a shows the average spectrum for all environmental conditions in the (0.01250)Hz(0.01-250)\,Hz range. From the calculated averaged spectra, it can be seen that:
@@
  • There is a clear distinction between the average spectrum for the noise and vessel conditions in the (2100)Hz(2-100)\,Hz, opening the path to derive discriminative features for ML tasks (see zoomed spectrum in Fig. 7.b).
@@
These conclusions guided our feature vector design, which will be composed of energy in logarithmically spaced frequency bands, to adequately represent lower frequency content. The considered bandwidth ranges from 4Hz4\,Hz to 98Hz98\,Hz, excluding the (4951)Hz(49-51)\,Hz interval to remove the 50Hz50\,Hz and 100Hz100\,Hz peaks shown in the spectral analysis (see Fig. 7.c), so that they will not mask the events captured by the DAS system.
Figure 9: Summary of frame processing strategies.
These results are consistent with the findings in previous literature [14], but we were not able to exploit frequency content that is known to be relevant in other studies, such as the 50Hz50\,Hz frequency band (in which we find a strong noise related harmonic) or that above 1kHz1\,kHz mentioned in [14], neither on the 100Hz200Hz100\,Hz-200\,Hz frequency band in [9], with a flat spectral response except for the harmonics described above. This lack of energy above 100Hz100\,Hz is in fact a direct consequence of the L=10mL=10\,m gauge length spatial sampling used in the interrogator setup. By the spatial Nyquist criterion an acoustic wavelength must be at least $2\timesgauge$ length (i.e. $\geq 20\,m$) to be resolved, and since the speed of sound in water is $\approx 1500\,m/s$, the upper resolvable frequency is $f_{\mathrm{max}}\,=\,\frac{c}{2L}\approx 75\,Hz.$ Any content above 75100Hz75–100\,Hz is therefore attenuated, which explains the flat response in the 100200Hz100–200\,Hz band, and, considering Figure 7.c, suggests that there is very relevant information in the 40100Hz40–100\,Hz frequency range.
As a prospective study of the feasibility of the proposed ML tasks, we generated a number of visualizations to relate the extracted features with the target elements, from which Fig. 8 shows an example. In the figure, the top graphic displays the temporal variation (horizontal axis, spanning for two hours) of the feature vector energy measured at each sensor (vertical axis, corresponding to $250$ sensors, spanning for 2553m2553\,m, from $14702$ to $17255$ meters in fiber distance from the interrogator). The bottom figure represents, at each time instant, the distance from the sensor range to the closest vessel, with the colorbar stating the color coding, between 0 and 2km2\,km (higher distances are saturated at the dark blue constant color to resemble lack of detected activity in the fiber). From Fig. 8, we can clearly see a very good match between the fiber measurements and the presence of vessels as they approach the fiber, cross over it, and progressively move away. We can also see different magnitudes of the vessel effect on the fiber (with longer or shorter spatial and temporal impacts), and with a clear relationship with the vessel presence in all the cases. Visible fiber response is detected along the full displayed fiber distance range, and up to roughly 10 minutes before and after the vessel crossing instant, which at an average speed of 10.2knots10.2\,knots implies a distance above 3km3\,km.
@@ VI-D Experimental Setup
We evaluated experimental conditions based on specific distance thresholds separating the vessel presence (Class 0) and absence (Class 1) for the classification task. The set of thresholds used was $D_{thr}\nobreak=\nobreak\{500,1000,1500,2000,3000,5000\}$ meters. For the vessel localization (regression) task, separate models were trained to estimate distances within ranges derived from these thresholds.
We also explored algorithmic strategies exploiting spatial and temporal redundancy:
  • Spatial redundancy: With configurations using $N_{S}=\nobreak\{10,25,50,100,250\}$ sensors, approximately spanning $100$, $250$, $500$, $1000$, and $2500$ meters, respectively.
  • Temporal redundancy: With temporal contexts of $S_{C}=\nobreak\{10,30,50\}$ seconds, using feature vectors from 1, 3, or 5 consecutive 10-second windows.
@@
Fig. 9 summarizes the described frame processing strategies.
@@
Figure 10: Baseline $F_{1}{}$-scores as a function of distance threshold (T[500,1000,1500,2000,3000,5000]_V1_S10_ACH), by varying the number of sensors used ($N_{S}=\nobreak\{10,25,50,100,250\}$): (a) $F_{1}^{G}$, (b) $F_{1}^{C0}$, (c) $F_{1}^{C1}$.
@@ VI-F Vessel Localization Experiments
A supplementary material Web page showing videos on the system operating in detection and distance estimation modes is linked from the project GitHub repository.
@@ VII Conclusions and Future Work
  • A data-driven spectral feature extraction based on logarithmically‐spaced frequency‐band energy measurements to capture vessel acoustic signatures.
@@
For the selected 1000m1000\,m threshold distance, our best performing vessel detection model achieved an overall $F_{1}{}$-score of over $90\%$, with balanced class-wise performance; and for vessel localization, the selected regression system achieved a mean absolute error of 141m141\,m, demonstrating reliable detection and distance estimation under diverse environmental and traffic conditions. These results highlight DAS as a practical alternative to conventional maritime surveillance methods, allowing for continuous monitoring without reliance on cooperative systems such as AIS.
The release of the full dataset used in this work will enable other researchers to easily reproduce our machine learning experiments, compare new algorithms, and extend vessel detection efforts in other submarine cable-related contexts.
Our field validation focused on vessel detection and localization for submarine cable protection represents the most extensive DAS-based vessel monitoring study to date, opening the path for several research areas to be explored as future work. Our first task will be devoted to applying the methodology to additional submarine cable deployments and emerging public DAS datasets [54], to assess generalization across different bathymetries, seabed couplings, vessel types, and traffic profiles. Beyond binary detection and range estimation, we will also address ML tasks to infer vessel speed, size, and type, as well as multi-target tracking and precise localization, by integrating richer AIS metadata, advanced signal processing and Deep Learning strategies. Exploiting meteorological data and bathymetric variables will be considered, either into feature sets or as model conditioning factors, given their significant influence on DAS submarine data acquisition. Finally, for real-world deployments, we will need to adopt efficient, low-latency implementations, designed for onboard DAS interrogators.

### v3

@@ V-B Spectral Feature Extraction
Extracting robust features from DAS data for vessel detection and localization is complicated due to the intricate physical processes converting ship noise into measured strain. Accurately modeling this conversion is challenging due to several factors: the complex and variable nature of noise radiated by different vessels [Rivet_2021]; the complexities of underwater acoustic propagation, including attenuation, reflection, and Doppler effects [Karasalo_2017, vadov2001long, li2024precise]; variations in the coupling between the fiber cable and seabed, which strongly influence strain measurements [Lior_2021, Mata_Flores_2023]; and potential artifacts introduced by the DAS system itself (due to electronic and optical design, and the conditions in the interrogator deployment environment).
Given that acquiring the precise knowledge needed for a full physics-based model is often impractical due to environmental variability and incomplete deployment information, our proposal is adopting a purely data-driven strategy. A frequent technique involves analyzing the energy within specific frequency bands, which has proven effective for interpreting DAS signals and providing features for associated ML tasks [tejedor2016toward, hartog2017introduction, Malaprade2019]. Adopting this principle, our work utilizes a data-driven spectral analysis, informed by available AIS data, to generate feature vectors suitable for guiding the ML process. Further details are given in Section VI-C, as it implies experimental work for feature vector design.
@@ V-C Machine Learning Tasks and Training Modules
  • XGBoost (eXtreme Gradient Boosting): An optimized gradient boosting framework that learns an ensemble of decision trees [Chen:2016:XST:2939672.2939785]. XGBoost was selected for its proven effectiveness, scalability to large datasets, and inherent mechanisms to mitigate overfitting, making it suitable for handling potentially high-dimensional feature spaces derived from DAS data.
@@
Another key element that impacts the reliability and robustness of a system and its real generalization capabilities is the rigor of the experimental approach [tejedor2017machine]. In time-series data, such as DAS recordings, temporal dependencies can introduce biases if not properly handled. The typically used random partitioning strategy may result in training and validation sets containing temporally adjacent data, leading to information leakage and overestimation of model performance. Special care must be taken in this stage, adopting specific cross-validation strategies that include restrictions on the selection of temporal ranges for training and testing subsets.
@@ V-D Result Post-Processing
In ML tasks dealing with time series, we can further improve the classification results by applying post-processing methods. Typical strategies are smoothing via sliding window averaging, median or mode filtering, majority voting (using ensembles or temporal ranges), multiple intervals, etc. [bagnall2017great].
Finally, one key benefit of DAS is the extensive range of sensed positions, which allows the systems to exploit the spatial and temporal redundancy of the acquired data. This should be explicitly included in either the feature extraction or ML stages. Suitable alternatives include exploiting array signal processing [munoz2022enhancing], or data integration across temporal and spatial ranges [tejedor2019contextual, huynh2025real], which will be detailed in Section VI-D.
@@ VI-A Database Partitioning
Given the characteristics of our task, we adopted the $k$-fold cross-validation strategy, a widely used technique for assessing model generalization capability [Wong2019-kfold]. This involves dividing the whole dataset into $k$ equal subsets (folds). The model is trained on k1k-1 folds and validated (tested) on the remaining fold. This process repeats $k$ times, with each fold serving once as the test set. The results from all test sets are then averaged to estimate the overall performance. Therefore, this method ensures that every data point is used for both training and testing, providing a comprehensive evaluation of the model’s performance while maintaining a strict separation between training and testing data.
@@ VI-B Performance Metrics
[…]
@@
In the vessel distance-estimation task, where the MAE will be calculated on data subsets with different distance thresholds, we will also normalize the MAE by these distance thresholds ($\delta_{thr}\in D_{thr}$), to get a relative estimation of the error, as compared with this distance threshold: $rMAE=\penalty 10000\frac{MAE}{\delta_{thr}}$.
To compare the results between the two ML approaches, we will quantify the uncertainty of our performance metrics using a non-parametric bootstrap procedure [efron1994introduction] (adapted to our $k$-fold cross-validation strategy) that has been further validated by subsequent works on bootstrap methods in ML [goo2024confidence, shabbir2024estimation].
@@ VI-C Feature Extraction Analysis
(a) Spectral content in the (0.1(0.1-250)Hz250)\,Hz range.
@@
Figure 7: Long term averaged strain power spectrum plots for strain measurements under different environmental conditions (expressed in dBs).
Figure 8: (a) Raw strain signal sensed at 15.4km15.4\,km in the time interval of the red dashed square in Fig. 8.d. (b) Filtered strain signal (4Hz4\,Hz to 98Hz98\,Hz, excluding the (4951)Hz(49-51)\,Hz) corresponding to the raw signal in Fig. 8.a and time-synchronized with it. (c) Filtered strain data as a function of time and distance along the cable corresponding to the red dashed square in Fig. 8.d, time synchronized with Figs. 8.a and 8.b. (d) Feature vector energy (in dBs) as a function of time and distance along the cable. (e) Distance from cable to nearest vessel (in meters), time-synchronized with Fig. 8.d.
Fig. 7.a shows the long term averaged strain power spectrum (expressed in dBs) for all environmental conditions in the (0.01250)Hz(0.01-250)\,Hz range. From the calculated averaged spectra, it can be seen that:
@@
  • There is a clear distinction between the average spectrum for the noise and vessel conditions in the (2100)Hz(2-100)\,Hz, opening the path to derive discriminative features for ML tasks (see zoomed spectra in Fig. 7.b).
@@
These conclusions guided our feature vector design, which will be composed of energy values in logarithmically spaced frequency bands (as shown in the logarithmic frequency axis in Fig. 7), to adequately represent lower frequency content. Each vector component is calculated by summing the FFT spectral coefficients corresponding to each defined frequency band. The considered bandwidth ranges from 4Hz4\,Hz to 98Hz98\,Hz, excluding the (4951)Hz(49-51)\,Hz interval to remove the 50Hz50\,Hz and 100Hz100\,Hz peaks shown in the spectral analysis (see Fig. 7.c), so that they will not mask the events captured by the DAS system. Fig. 8.b shows an example of the filtered strain signal corresponding to the raw strain signal in Fig. 8.a. An example of the evolution of the filtered strain as a function of time and distance along the cable is shown in Fig. 8.c.
[…]
As a prospective study of the feasibility of the proposed ML tasks, we generated a number of visualizations to relate the extracted features with the target elements, from which Figs. 8.d+e show an example. Fig. 8.d displays the temporal variation (horizontal axis, spanning for two hours) of the feature vector energy measured at each sensor (vertical axis, corresponding to $250$ sensors, spanning for 2553m2553\,m, from $14702$ to $17255$ meters in fiber distance from the interrogator). The feature vector energy is calculated as the sum of the energy components in all the selected frequency bands, and getting its value in db. Fig. 8.e is time synchronized with Fig. 8.d and represents, at each time instant, the distance from the sensor range to the closest vessel, between 0 and 2km2\,km (higher distances are saturated at the dark blue constant color to resemble lack of detected activity in the fiber). From Figs. 8.d+e, we can clearly see a very good match between the fiber measurements and the presence of vessels as they approach the fiber, cross over it, and progressively move away. We can also see different magnitudes of the vessel effect on the fiber (with longer or shorter spatial and temporal impacts), and with a clear relationship with the vessel presence in all the cases. Visible fiber response is detected along the full displayed fiber distance range, and up to roughly 10 minutes before and after the vessel crossing instant, which at an average speed of 10.2knots10.2\,knots implies a distance above 3km3\,km.
@@ VI-D Experimental Setup
We evaluated experimental conditions based on specific distance thresholds separating the vessel presence (Class 0) and absence (Class 1) for the classification task. The set of thresholds used, measured in meters, was $D_{thr}\penalty 10000=\penalty 10000\{500,1000,1500,2000,3000,5000\}$. For the vessel localization (regression) task, separate models were trained to estimate distances within ranges derived from these thresholds.
We also explored algorithmic strategies exploiting spatial and temporal redundancy (graphically summarized in Fig. 9):
  • Spatial redundancy: With configurations using $N_{S}=\penalty 10000\{10,25,50,100,250\}$ sensors, approximately spanning $100$, $250$, $500$, $1000$, and $2500$ meters, respectively.
  • Temporal redundancy: With temporal contexts of $S_{C}=\penalty 10000\{10,30,50\}$ seconds, using feature vectors from 1, 3, or 5 consecutive 10-second windows.
@@
Figure 9: Summary of frame processing strategies to exploit spatial and temporal redundancy in the feature space, and majority voting in the output results.
@@
Figure 10: Baseline $F_{1}{}$-scores as a function of distance threshold (T[500,1000,1500,2000,3000,5000]_V1_S10_ACH), by varying the number of sensors used ($N_{S}=\penalty 10000\{10,25,50,100,250\}$): (a) $F_{1}^{G}$, (b) $F_{1}^{C0}$, (c) $F_{1}^{C1}$.
@@ VI-F Vessel Localization Experiments
A supplementary material Web page showing videos on the system operating in detection and distance estimation modes is available at https://geintra-uah.org/psi/index.html and is also linked from the project GitHub repository.
@@
VI-G Comparison with beamforming methods
To provide a reference baseline based on classical acoustic localization techniques, we implemented a beamforming-based distance estimation method following the approach presented by Paap et al. [paap2025leveragin], who conducted their study in the same geographical area as ours. In particular, we adopted the Steered Response Power (SRP) algorithm [dibiase2001robust], which is virtually identical to the implementation of the migration-based method used by Paap (sharing the same propagation assumptions and processing principles), but calculating power in each grid position instead of stacked absolute values of the DAS waveforms. The SRP algorithm was applied to the DAS strain data using the same 250 channels and 10-second temporal windows than in our experimental setup. To optimize the propagation model parameters, we applied Paap’s methodology [paap2025leveragin], scanning for the optimal sound speed by maximizing the coherency of back-projected waveforms. We tested speeds from 100m/s100\,m/s to 1900m/s1900\,m/s and determined a maximum response at 1750m/s1750\,m/s. This value is highly consistent with the 1700m/s1700\,m/s found by Paap et al. in the same region [paap2025leveragin] (also with a buried cable), and can be interpreted as evidence of significant acoustic energy transmission through the seabed top sediment layer, rather than through the water column.
For each time window, the SRP function was evaluated over a two-dimensional spatial grid with 20m20\,m resolution covering the vicinity of the monitored cable (a bounding box defined as 1000m1000\,m from the extreme cable positions of the used channels). The algorithm determines the most probable vessel position as the grid point maximizing the steered response power, from which we derived the corresponding distance to the cable. All results presented in this section correspond to vessels located within 1km1\,km of the cable.
Figure 19: Comparison of SRP beamforming and ML regression methods. (a) SRP map for a case where both methods perform well, particularly as the vessel is close to the cable. (b) SRP map for a vessel that is more difficult to detect; while both methods struggle, the ML regressor’s error is lower. (c) A time-series plot of a vessel crossing above the fiber, showing the distance estimations by SRP (red trace) and the ML regressor (gray trace) against the ground truth (green trace), illustrating a case where both methods track the vessel accurately. (d) A distance estimation time-series plot for a quieter or more challenging vessel, demonstrating a case where the ML regressor performs significantly better than the SRP method.
TABLE II: Distance estimation performance of SRP and ML-based regression for vessels within 1 km of the cable.
Method MAE [m] 95 % CI [m]
SRP Beamforming 225 [222.17, 227.97]
ML Regression (XGBoost) 171 [169.22, 172.83]
Table II summarizes the MAE results with the 95 % confidence intervals (CI) for both methods. Fig. 19.a shows a sample SRP map where the steered-response peak accurately identifies the vessel position, a scenario where both the beamformer and the ML regressor perform very well. In contrast, 19.b shows a more challenging case where the vessel is harder for both methods to detect, though the ML regressor ultimately yields a lower estimation error. The time-series plots in Figs. 19.c and 19.d compare the distance estimations over a time window centered on a vessel crossing, showing the true distance (estimated from AIS), the ML regressor prediction, and the SRP prediction. Fig. 19.c shows an example where both methods track the vessel effectively, whereas Fig. 19.d shows an example with a more challenging vessel where the ML regressor performs much better than the beamforming approach due to the presence of local maxima in the SRP map that do not correspond to the real vessel position.
Beamforming methods provide physically interpretable spatial maps, require no training data, and adapt easily to new deployments. However, their performance is highly sensitive to propagation model parameters that are difficult to control, such as the effective sound speed and the fiber-seabed coupling. Additionally, they require dense spatial grids that makes them computationally expensive, especially in long-cable deployments.
Figure 20: Distribution of absolute distance errors by (a) vessel size range, (b) vessel type (indicating also the average length (avg.len.) for the group type).
In contrast, the proposed ML-based approach avoids strong physical assumptions on the propagation model but depend on representative training data. Once trained, models validated on diverse cable segments can generalize well to new environments, providing faster and more accurate distance estimates by directly mapping signal features to vessel range, effectively exploiting spatiotemporal redundancies across sensors, and offering a more scalable and practical framework for continuous vessel monitoring.
VI-H Dependency on Vessel Size and Type
To assess the influence of vessel characteristics on localization accuracy, we analyzed the absolute distance error as a function of both vessel size and type. We evaluated the task with the 1000m1000\,m distance threshold, 250 channels and 10-second temporal windows (that achieved a MAE of 171m171\,m), with the results distribution boxplots shown in Fig. 20.
In general, larger vessels show the lowest errors, consistent with their expected stronger and more stable acoustic emissions. However, there are cases not fitting in this tendency, such as the Passenger Ships (mainly cruises), achieving higher error than the smaller Pilot Vessels, probably due to their expected lower radiated noise (being focused in passenger comfort).
As a conclusion, vessel size and type affect DAS-based localization performance, and its detailed study is open for future work, initially suggesting that class dependent models could further enhance system robustness and enable addressing vessel type identification tasks.
@@ VII Conclusions and Future Work
  • A data-driven spectral feature extraction based on logarithmically-spaced frequency‐band energy measurements to capture vessel acoustic signatures.
@@
For the selected 1000m1000\,m threshold distance, our best performing vessel detection model achieved an overall $F_{1}{}$-score of over $90\%$, with balanced class-wise performance; and for vessel localization, the selected regression system achieved a mean absolute error of 141m141\,m, demonstrating reliable detection and distance estimation under diverse environmental and traffic conditions. We also carried out a comparison with a beamforming-based distance estimation method, showing that the proposed ML approach achieves lower localization errors and offers a more robust and computationally efficient alternative. These results highlight DAS as a practical alternative to conventional maritime surveillance methods, allowing for continuous, all-weather monitoring without reliance on cooperative systems such as AIS.
The release of the full dataset used in this work will enable other researchers to easily reproduce our machine learning experiments, compare new algorithms, and extend vessel detection efforts in other submarine cable-related contexts. Our field validation focused on vessel detection and localization for submarine cable protection represents the most extensive DAS-based vessel monitoring study to date, opening the path for several research areas to be explored as future work. Our first task will be devoted to applying the methodology to additional submarine cable deployments and emerging public DAS datasets [ais-data-ieee-dataport-1308-8605-25], to assess generalization across different bathymetries, seabed couplings, vessel types (proved to have an impact in localization accuracy), traffic profiles, and weather conditions. Beyond binary detection and range estimation, we will also address ML tasks to infer vessel speed, size, and type, as well as multi-target tracking and precise localization, by integrating richer AIS metadata, advanced signal processing and Deep Learning strategies. Exploiting meteorological data and bathymetric variables will be considered, either into feature sets or as model conditioning factors, given their significant influence on DAS submarine data acquisition. Finally, for real-world deployments, we will need to adopt efficient, low-latency monitoring pipelines such as the recent proposal in [biondi2025real].